OpenStudy (anonymous):

Find the slope of the line that passes through the points (-1, 3) and (3, 5). A. 1/2 B. 1 C.2/5 D.2 When i got the test back i circled B and now that i did it i got A so i want to make sure if its A

OpenStudy (anonymous):

\[m=\frac{ (y2-y1) }{ (x2-x1) }\]

hartnn (hartnn):

yes, its A

OpenStudy (anonymous):

How did you get A ?

hartnn (hartnn):

m= (5-3) / (3-(-1)) = 2/(3+1) = 2/4 = 1/2

hartnn (hartnn):

got that ^ ?

OpenStudy (anonymous):

Thanks!

hartnn (hartnn):

welcome ^_^